If 149.1 g of KCl is dissolved in water to make 500 mL solution, what is the molarity of the solution?


1
Expert's answer
2021-02-18T07:29:12-0500

n(KCl)=m(KCl)/Mr(KCl)=149.1/74.5= 2 mole

Molarity= 2 mole/ 0.5 l= 4 mole/L